OpenStudy (anonymous):

To find the height of a tall tree, a surveyor moves 140 feet away from the base of the tree and then, with a transit 4 feet tall, measures the angle of elevation to the top of the tree to be 53°. What is the height of the tree? Round your answer to the nearest foot. A. 190 ft B. 186 ft C. 105 ft D. 109 ft

OpenStudy (anonymous):

@Marco,Phillip

OpenStudy (anonymous):

A. 190 ft. Form a right angled triangle from the following points: the transit, a point 4 feet up the tree and the top of the tree. So since tan = opp/adj, in this case tan(53) = (h-4)/140, where h is the height of the tree and (h-4) is the distance between the top of the tree and the point 4 feet up the tree that we chose to form a right angled triangle. Thus h-4 = 140 * tan(53) = 186 and so h = 190 feet.

OpenStudy (anonymous):

thx XD!!! I have one more question, will you help me with it?

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

it's true and false

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The length of side x could most easily be found by using the tangent of angle a.

OpenStudy (anonymous):

I have to draw

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):



OpenStudy (anonymous):

from what i see on other i would say it is true

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem